A vehicle purchased for $6,400 decreases in value at a rate of 9.3% per year. After how many years will the vehicle be worth $3,800? Round your answers to the nearest tenth of a year.

Answer by Cromlix(4381)

Hi there,
100% - 9.3% = 90.7%
90.7% = 90.7/100 = 0.907
$6,400 x 0.907^x = $3,800
0.907^x = $3,800/$6,400
0.907^x = 0.59375
x log(0.907) = log (0.59375)
x = log(0.59375)/ log(0.907)
x = 5.34 years
x = 5.3 years (nearest tenth of a year)
Hope this helps :-)